Grade 11Physical Chemistry

please clear the calculation means explain it

Number of moles of N2 = 50 kg of N2 = 50 X 10 3 g/1 kg x 28g = 17.86 x 10 2 mole

Number of moles of H2 = 10 kg of N2 = 10 X 103 g/ 1 kg x 2 = 4.96X 103 mol

prachi
8 Years agoGrade 11

1 Answer

Adarsh
8 Years ago
Dear,Here we have to calculate no more moles of N2 in 50 kg of it .28g of dinitrogen have 1 mole molecule in it , so 50000g of nitrogen has 50000/28 mole = 17.86 X 10^2.Now no of mole of H2 in 10 kg H2 = 10000/2=5000mol Which is nearly equal to 4.96 X 10^ 3 mole.Hope that this will help you ..RegardsAdarsh